ANALYSIS OF THE HYPERFINE STRUCTURE IN THE [20.6]5 - X4 TRANSITION OF COBALT MON-FLUORIDE.

\underline{A. G. ADAM} and WEI SHA, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

The high resolution visible laser spectrum of the 484 nm band system of cobalt monofluoride has been obtained. The molecules were produced by reaction of laser ablated cobalt atoms with 1% SF_{6} seeded in He under supersonic jet conditions. The 0-0 and 1-0 bands of the system were probed with a Coherent 699-29 cw ring dye laser giving experimental linewidths on the order of 200 MHz. The appearance of the spectrum at low resolution where a strong R-, and weaker Q-, and P-branches were observed is indicative of ΔΛ = ΔΩ =+1. This and analysis of the first lines yields a value of omega=4 for the ground state, consistent with earlier work, and a value of omega=5 for the upper state. Comparison with calculations on CoH has allowed this transition to be assigned as the [20.6]^{3}Γ_{5} - X^{3}Φ_{4} transition. The molecular constants derived from a nonlinear least squares fit to the data will be presented.